OpenStudy (anonymous):

What is the x-intercept of the line with the equation ? 1/3x+y=-15

OpenStudy (jonnyvonny):

x intercepts happen when y=0. First off, get y alone to get y=-1/3x-15 Set y equal to 0 0=-1/3x-15 Now get x alone. Tell me what you get, and I'll post what I got.

OpenStudy (jonnyvonny):

x=-45, I don't have time for someone who doesn't care.